I x+4.2I + 3.6=16.2
Answer by stanbon(75887)

I x+4.2I + 3.6=16.2
-------------------------
|x+4.2| = 12.6
---
x + 4.2 = 12.6 or x+4.2 = -12.6
x = 8.4 or x = -16.8
===============================
Cheers,
stan H.